OpenStudy (anonymous):

rewrite the following linear equation in slope-intercept form. Write your answer with no spaces y+4=-2(x-1)

OpenStudy (anonymous):

@pooja195 ??

pooja195 (pooja195):

y+4=-2(x-1) Start by distributing the 2 into the parentheses

pooja195 (pooja195):

*-2

pooja195 (pooja195):

-2 x X=?

OpenStudy (anonymous):

-2x*-2 @pooja195

pooja195 (pooja195):

Good Now we have y+4=-2x-2 Subtract 4 from both sides what do you have?

OpenStudy (anonymous):

y=-2x-6?

pooja195 (pooja195):

Well done :) \(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)

OpenStudy (anonymous):

I might have a few more questions later on if thats fine with you @pooja195

pooja195 (pooja195):

Just tag me =)